OpenStudy (anonymous):

Expand and evaluate the power. The third power of 2.5

OpenStudy (anonymous):

the third power of 2.5 would be 2.5*2.5*2.5 so 2.5 *2.5 = 6.25 6.25 * 2.5 = 15.625 \[2.5^3 = 15.625\]

OpenStudy (anonymous):

thanks